how much does an alloy that is 20% copper should be mixed with 200 onces of an alloy that is 90% copper in order to get an alloy that is 30% copper
Answer by stanbon(75887)

how much does an alloy that is 20% copper should be mixed with 200 onces of an alloy that is 90% copper in order to get an alloy that is 30% copper
-----------------
Equation:
copper + copper = copper
0.20x + 0.90*200 = 0.30(x+200)
Multiply thru by 100 to get:
20x + 90*200 = 30x + 30*200
-10x = -60*200
x = 1200 onces (amt of 20% alloy needed in the mixture)
==========================================================
Cheers,
Stan H.